# Openpi Comet: Competition Solution For 2025 BEHAVIOR Challenge

### v1

@@ Abstract
The 2025 BEHAVIOR Challenge is designed to rigorously track progress toward solving long-horizon tasks by physical agents in simulated environments. BEHAVIOR-1K focuses on everyday household tasks that people most want robots to assist with and these tasks introduce long-horizon mobile manipulation challenges in realistic settings, bridging the gap between current research and real-world, human-centric applications. This report presents our solution to the 2025 BEHAVIOR Challenge in a very close 2nd place and substantially outperforms the rest of the submissions. Building on $\pi_{0.5}$, we focus on systematically building our solution by studying the effects of training techniques and data. Through careful ablations, we show the scaling power in pre-training and post-training phases for competitive performance. We summarize our practical lessons and design recommendations that we hope will provide actionable insights for the broader embodied AI community when adapting powerful foundation models to complex embodied scenarios.
@@ 1 Introduction
In this report, we examine how far a strong publicly available VLA backbone can be pushed on long-horizon tasks using careful data, training, and inference design within a simple end-to-end training pipeline. We treat the BEHAVIOR Challenge as a case study in adapting powerful but generic foundation policies to a complex embodied benchmark. Through systematic exploration of training configurations, pre-training choices, and inference strategies, we show that our solution completes 22 tasks out of the 50 household tasks, achieving a Q-score of 0.2514 in the competition Table 1.
@@ 2 Architecture
Table 1: Results of 2025 BEHAVIOR Challenge for standard track. Q-score for the test set (bold) is used for final ranking.
@@
2 Comet (ours) 0.1440 0.1140 0.1830111Due to limited time by the submission deadline, we could not finish evaluation across all tasks on time. Thus, our submitted validation Q-score is lower than the actual validation Q-score reported in section 4. 0.2514
@@ 4 Experiment
Pre-trainin Post-training Theoretical Best
0.19 0.22 0.31
@@ 4.2 Pre-training
After pre-training, we reach a validation Q-score of 0.19 on the validation set, as shown in section 4.
@@ 4.3 Post-training
Despite the recent progress in online RL (chen2025pi_), its low sample efficiency makes it impractical for the BEHAVIOR challenge. Moreover, online RL requires a heterogeneous compute setup: GPUs with RT Cores are needed for simulation, while GPUs with Tensor Cores are needed for model training and rollouts. Given these constraints, we instead adopt rejection sampling fine-tuning (RFT), a technique shown to be effective in both LLMs and VLMs (ahn2024large; touvron2023llama; azzolini2025cosmos). An overview of our RFT setup is provided in Algorithm 1. Starting from all the provided train and validation human demonstrations for each scene, we randomly perturb the robot’s initial pose and use our pre-trained policy to perform rollouts under these perturbed configurations. Using both train and validation set helps avoid overfitting to the validation set and provides better signals of the model quality. Successful rollouts are retained as additional demonstrations. We perform $N=3$ rounds of RFT in total, collecting on average $T=8500$ trajectories per round, and eventually selected approximately 2500 trajectories for training after de-duplication and task balancing.
@@
After post-training, our suite of models achieves a validation Q-score of 0.22. In addition to the effectiveness, RFT provides an estimate of the theoretical upper bound on model performance by computing the union of success instances across all past model checkpoints. As shown in Figure 4 and section 4, the theoretical best validation Q-score we could achieve is 0.31. Due to time constraints, we were unable to fully close this gap, leaving additional headroom for further post-training. Yet, the theoretical best Q-score provides a tangible target of additional improvements. Please refer to Figure 5 for representative visualizations of policy executions results.
Figure 4: By selecting the best post-trained policy model per task instance, we can achieve an aggregated validation Q-score of 0.31, serving as the upper bound of our model's performance.
@@ 4.4 Ablations
#1 Temporal Ensemble 50 RGB
Head:224
Wrist:224
0.00
@@
#2 Receding Horizon 8 RGB
Head:224
Wrist:224
0.00
@@
#3 Receding Horizon 32 RGB+Depth Image
Head:224
Wrist:224
0.20
@@
#4 Receding Horizon 32 RGB
Head:224
Wrist:224
0.30
Head:720
Wrist:480
0.60

### v3

@@ Abstract
The 2025 BEHAVIOR Challenge is designed to rigorously track progress toward solving long-horizon tasks by physical agents in simulated environments. BEHAVIOR-1K focuses on everyday household tasks that people most want robots to assist with and these tasks introduce long-horizon mobile manipulation challenges in realistic settings, bridging the gap between current research and real-world, human-centric applications. This report presents our solution to the 2025 BEHAVIOR Challenge in a very close 2nd place and substantially outperforms the rest of the submissions. Building on $\pi_{0.5}$, we focus on systematically building our solution by studying the effects of training techniques and data. Through careful ablation studies, we reveal the scaling benefits in both the pre-training and post-training phases, leading to a validation Q-score of 0.345, significantly surpassing previous state-of-the-art performance. We summarize our practical lessons and design recommendations that we hope will provide actionable insights for the broader embodied AI community when adapting powerful foundation models to complex embodied scenarios.
@@ 1 Introduction
In this report, we examine how far a strong publicly available VLA backbone can be pushed on long-horizon tasks using careful data, training, and inference design within a simple end-to-end training pipeline. We treat the BEHAVIOR Challenge as a case study in adapting powerful but generic foundation policies to a complex embodied benchmark. Through systematic exploration of training configurations, pre-training choices, and inference strategies, we show that our solution completes 22 tasks out of the 50 household tasks, achieving a Q-score of 0.2514 in the competition Table 1. After the challenge, we have refined our post-training algorithm and achieved a validation Q-score of 0.3453, substantially outperforming all previous results.
@@ 2 Architecture
Table 1: Results of 2025 BEHAVIOR Challenge for standard track and our post-challenge solution. Q-score for the test set is used for final ranking. Best viewd in color.
@@
2 Comet (ours) 0.1440 0.1140 0.1830111Due to limited time by the submission deadline, we could not finish evaluation across all tasks. 0.2514
@@ 4 Experiment
Pre-training Post-training Theoretical best
0.192 0.345 0.611
@@ 4.2 Pre-training
After pre-training, we reach a validation Q-score of 0.192 on the validation set, as shown in section 4.
@@ 4.3 Post-training
Despite the recent progress in online RL (chen2025pi_), its low sample efficiency makes it impractical for the BEHAVIOR challenge. Moreover, online RL requires a heterogeneous compute setup: GPUs with RT Cores are needed for simulation, while GPUs with Tensor Cores are needed for model training and rollouts. Given these constraints, we instead adopt rejection sampling fine-tuning (RFT), a technique shown to be effective in both LLMs and VLMs (ahn2024large; touvron2023llama; azzolini2025cosmos). An overview of our RFT setup is provided in Algorithm 1. Starting from all the provided train and validation human demonstrations for each scene, we randomly perturb the robot’s initial pose and use our pre-trained policy to perform rollouts under these perturbed configurations. Using both train and validation set helps avoid overfitting to the validation set and provides better signals of the model quality. Successful rollouts are retained as additional demonstrations. We perform $N=3$ rounds of RFT in total, collecting on average $T=8500$ trajectories per round, and eventually selected 1469 trajectories for training after de-duplication and task balancing.
@@
During the challenge, our suite of models achieves a validation Q-score of 0.224 after post-training. After the challenge, we further refine the task balancing strategy and achieve a significantly higher Q-score of 0.345 on the validation set, using only two model checkpoints. Representative policy execution visualizations are provided in Figure 5. Beyond improving overall performance, Rejection Sampling Fine-Tuning (RFT) also serves as a diagnostic mechanism by aggregating successful instances across historical checkpoints. As shown in section 4, the theoretical best results we could obtain is a validation Q-score of 0.611 and a success rate of 0.35, revealing substantial headroom for further optimization. Further scaling up the model size could be helpful to realize the full potential of our post-training strategy.
Figure 4: We can achieve an aggregated validation Q-score of 0.345 and a validation success rate of 15% across 50 tasks. Our models can perform easy tasks such as turnining_on_radio robustly. Zero-success rate tasks are not shown.
@@ 4.4 Ablations
#1 Temporal Ensemble 50 RGB Head:224 Wrist:224 0.00
@@
#2 Receding Horizon 8 RGB Head:224 Wrist:224 0.00
@@
#3 Receding Horizon 32 RGB+Depth Image Head:224 Wrist:224 0.20
@@
#4 Receding Horizon 32 RGB Head:224 Wrist:224 0.30
Head:720 Wrist:480 0.60